### v1

# Cutting through buggy adversarial example defenses: fixing 1 line of code breaks Sabre

Abstract
Sabre is a defense to adversarial examples that was accepted at IEEE S&P 2024. We first reveal significant flaws in the evaluation that point to clear signs of gradient masking. We then show the cause of this gradient masking: a bug in the original evaluation code. By fixing a single line of code in the original repository, we reduce Sabre’s robust accuracy to 0%. In response to this, the authors modify the defense and introduce a new defense component not described in the original paper. But this fix contains a second bug; modifying one more line of code reduces robust accuracy to below baseline levels.
1 Introduction
Sabre [5] is a defense to adversarial examples that claims to be 3×3$\times$3 × more robust to attack than the current state-of-the-art [robustbench], bringing the best attacks on CIFAR-10 at a distortion of $\varepsilon=8/255$ from 29% success rate down to just 11%. In this paper we begin with a critique of the Sabre paper, and show it contains multiple evaluation flaws. For example, the paper claims that Sabre is more accurate when an adversary is attacking the defense than when not under attack, claims mathematically impossible levels of robustness, and does not follow standard recommended best practices [3] when evaluating adversarial robustness—including omitting any analysis against adaptive attacks [7].
Nevertheless, Sabre was accepted at IEEE S&P 2024, one of the top computer security conferences. (Readers may be forgiven for mistaking this defense for a different defense—also with a flawed evaluation—accepted at last year’s IEEE S&P 2023. [2])
In this short paper we show that by changing a single line of code to fix a bug in the published evaluation code, the attack success rate increases to 100%. The authors responded to our attack by modifying the published defense code to introduce a new defense component that was not present in the original paper, thereby changing the defense to a different defense than was initially proposed. Unfortunately, this modification contains a second bug, and by modifying a second line of code we again completely break the defense. The authors claim to have had a fix to this second attack for six weeks, but we have been unable to verify these claims because the authors have not published code for this fix, nor have they described the necessary changes. When the authors provide this information we will update our paper with an analysis of any future modifications or corrections to the currently published defense implementation.
2 Background & Sabre
We assume familiarity with evaluating the robustness of adversarial example defenses [3]. We encourage readers not familiar with this literature to read the related work before returning to this paper.
Sabre [5] is a preprocessor defense: given an input $x$, it first pre-processes the input with a function $p$, and then classifies the resulting pre-processed image with a standard (adversarially trained) neural network $f$; the resulting defended classifier is thus $f_{\textsc{Sabre}}=f\circ p$. Because our attack is completely independent of the method by which Sabre works, we do not describe it in any more detail. Interested readers can refer to the original paper [5] for a complete description of the methodology.
3 Mathematically Impossible
Sabre makes a number of claims that are impossible for any correct evaluation. The impossibility of the claims it makes are sufficiently well understood that, instead of describing the errors ourselves, we prompt GPT-4 with a copy of the checklist in [3], a screenshot of the defense paper [5], and then ask the question “Give one thing that is concerning about this image that suggests the paper’s evaluation might be flawed. Focus on [some paragraph or figure].”
  • Nontrivial accuracy at $\varepsilon=0.5$. “Given the estimated value of approximately 92.66% accuracy at a perturbation bound ($\varepsilon$) of 0.5 for the.“Ours” defense, this result appears to contradict the expectations set forth by Carlini et al. They stated that for an $\ell_{\infty}$ distortion of 0.5, model accuracy should not be better than random guessing because any image can be converted into a solid gray picture, making it impossible for the model to accurately predict the original content.”
  • Attacking the model makes it more accurate. “In the SPEECH-COMMANDS row evaluated with SABRE (the proposed defense), there is a noticeable discrepancy: the performance under the “No Attack” condition is 83.57, which is lower than the performance under several of the adversarial attacks such as FGSM (91.21), PGD 7 steps (95.40), and PGD 20 steps (95.68). This is concerning because we would generally expect a model to perform best in the absence of any attack, as adversarial examples are specifically designed to degrade performance.
    The fact that the model performs better under attack than when there is no attack challenges the typical expectations and could suggest issues such as overfitting to the adversarial examples, a problem with the evaluation methodology, or that the reported numbers may not accurately reflect the true robustness of the model. This anomaly warrants a closer examination of the experimental setup and the robustness evaluation to ensure the validity of the results.”
In both of these cases, the language model has correctly identified the flaws in the original paper, and the author of this paper has checked that the model’s response is accurate and true.
4 Additional evaluation flaws
Sabre also deviates from recommended practices in evaluating adversarial robustness. Specifically, it does not follow the widely accepted best practices in evaluating adversarial robustness [3] in the following ways:
  • [3](§5.2) “Verify iterative attacks perform better than single-step attacks” In the original evaluation, the robust accuracy at 100 iterations of PGD is higher than the robust accuracy at 1 step of FGSM.
  • [3](§5.4) “Verify adaptive attacks perform better than any other” The paper performs no adaptive attack. (See the following section.)
  • [3](§4.4) “Try at least one gradient-free attack and one hard-label attack” The paper performs no gradient-free attacks or hard-label attack.
  • [3](§5.1) “Compare against prior work and explain important differences” The paper includes a comparison to PGD, but vastly under-reports the robustness of PGD as 30% accuracy at $\varepsilon=0.3$ instead of 90%+. (It also omits comparisons to Feature Squeezing; see the following section.)
  • [3](§4.11) “Attack with random noise of the correct norm” The paper does not attempt this evaluation.
  • [3](§4.13) “Perform ablation studies with combinations of defense components removed” The paper does not do this. (See the following section.)
  • [3](§4.3) “Applying many nearly-identical attacks is not useful” The paper reports the robustness of multiple nearly-identical attacks instead of performing strong adaptive attacks.
5 Significant flaws not observable from the paper
We make a number of other comments on the Sabre defense that can not be observed by reading the paper alone, and can only be noticed by reading the published code or corresponding with the authors.
Sabre is not evaluated against adaptive attacks.
An adaptive attack is one that is “adapted to the specific details of the defense and attempt[s] to invalidate the robustness claims that are made” [3] “Adaptive attacks have (rightfully) become the de facto standard for evaluating defenses to adversarial examples.” [7] As the authors of [3] say, “Defending against non-adaptive attacks is necessary but not sufficient. It is our firm belief that an evaluation against non-adaptive attacks is of very limited utility” (emphasis not added here, and was present in original text).
Despite the fact that the paper writes Sabre “is virtually as robust against adaptive attacks as it is against weak attacks,” this was not actually evaluated. The authors do not attempt an adaptive attack of Sabre, and in personal communication, the authors state that Sabre is not intended to be robust against future attacks—only those specific attacks considered in the paper. When the paper writes it is robust against adaptive attacks, the author’s intent was to say that that AutoAttack [4], one particular adversarial example generation method, does not succeed more often than PGD.
But this is not what an adaptive attack means. No attack can be said to be adaptive isolation—indeed, the phrase “adaptive attack” never appears in the AutoAttack paper [4]. AutoAttack was not meant to be a replacement for an adaptive attack where the adversary designs the attack to specifically target the defense: they even write “we […] see AutoAttack and adaptive attacks as complementary. Claiming robustness to adaptive attackers without actually performing an adaptive attack misleads the reader.
Sabre includes baselines against known-insecure defenses.
The paper compares the proposed defense to ME-Net [9], a defense published in December 2019 that was broken two months later in February 2020 [7]. ME-Net is indeed a similar defense, in that it also pre-processes inputs in a similar manner to Sabre, and so one might expect that Sabre would evaluate against a similar attack strategy as the one that was used to break ME-Net, but this was not done.
Moreover, the authors state that they were aware that ME-Net was previously broken, and yet (1) did not mention this fact in their paper, and (2) did not attempt to break their defense with similar attacks to the ones that broke ME-Net. Papers should not compare to prior insecure defenses without considering the attack that broke them, because it is likely that the attack that broke the prior paper also will break this one. And as we will show, the type of gradient obfuscation ME-Net suffered from is very similar to the type of gradient obfuscation Sabre suffers from.
Sabre implements prior baselines incorrectly.
The Sabre paper contains a comparison to adversarial training [6], the leading methodology behind the current state-of-the-art adversarial example defenses. Unfortunately, Sabre incorrectly implements adversarial training. On MNIST, instead of adversarial training reducing attack success rate to just 11% on at a distortion of $\varepsilon=0.3$, the paper reports that the attack success rate against adversarial training is over 70%. This makes Sabre appear to be a much larger improvement compared to prior work than is actually the case. While qualified reviewers should know the correct baseline numbers to which Sabre should be compared, an under-informed reader not familiar with prior work might (incorrectly) believe that Sabre has improved significantly on prior work when, in fact, the baseline is instead incorrect.
(Sabre also omits any comparisons to defenses that have since improved on adversarial training since it was published in 2018; further reducing the gap between the papers state-of-the-art MNIST claims and the true state-of-the-art.)
6 Our attack
We break Sabre twice.
6.1 Our first break
The paper repeatedly reports that Sabre is completely differentiable, but the defense is evaluated by wrapping the preprocessing step with BPDA [1], a method that replaces the gradient of the preprocessor with an estimate that $f^{\prime}(x)=x$. But Sabre is already differentiable; wrapping it with BDPA removes the useful gradient and replaces it with one that is much less useful. And so to break Sabre all we have to do is remove the unnecessary BPDA wrapping.
1diff --git a/core/defenses/sabre.py b/core/defenses/sabre.py
2index fe509e6..bf13629 100644
3--- a/core/defenses/sabre.py
4$+$ b/core/defenses/sabre.py
5@@ -165,7 +165,7 @@ class SabreWrapper(nn.Module):
6 model = Sabre(eps=eps, wave=wave, use_rand=use_rand, n_variants=n_variants)
7 self.core = model
8 self.base_model = base_model
9- self.transform = BPDAWrapper(lambda x, lambda_r: model.transform(x, lambda_r).float())
10+ self.transform = (lambda x, lambda_r: model.transform(x, lambda_r).float())
11 @property
12 def lambda_r(self):
With this change, after running the original evaluation code, the defense achieves $<$5% accuracy on the MNIST dataset with a perturbation bound of 0.3. By increasing the number of attack iterations from 7 to 100, the defense accuracy drops to 0%. On the CIFAR-10 dataset we can achieve 0% accuracy at a perturbation bound of 8/255 (as originally considered in the paper) with the same bugfix.
6.2 A fix to our first break
The authors acknowledge that our break effectively reduces Sabre’s robust accuracy to 0%. In response, the authors modify the defense and introduce a new component that is not described in the original paper. Specifically, the modified Sabre first discretizes inputs to a fixed number of digits of precision as follows:
1def precision_blend(self, x):
2 if self.eps > 0:
3 precision = max(min(-int(math.floor(math.log10(abs(1.25 * self.eps)))) - 1, 1), 0)
4 x = self.diff_round(x, decimals=precision)
5 return x
6
7def diff_round(self, x, decimals=1):
8 scale_factor = (10 ** decimals)
9 x = x * scale_factor
10 diff = (1 + self.error_coefficient) * x - torch.floor(x)
11 x = x - diff + torch.where(diff >= 0.5, 1, 0)
12 x = x / scale_factor
13 return x
Nevertheless, the authors claim that this defense component was intended as part of the original defense. In personal communication with the authors, they assert this defense component was not described in the paper because it is not an important detail of the overall defense. (This is despite the fact that the entire reason this component was introduced in the first place was to prevent the attack we described above.)
In further communication with the authors, they informed us that hyperparmaters should be set to discretize MNIST images so that every pixel is set to either 0 or 1 (i.e., zero decimals of precision in the code above), and discretize CIFAR-10 images to one decimal digit of precision.
On differentiability.
Throughout the paper, Sabre claims to be “end-to-end differentiable to avoid gradient masking”, claims “innate differentiability of [their] defense mechanism”, and claims to “ensure end-to-end differentiability” [5]. But with this new changed proposed by the authors, this is no longer the case: discretization is the quinticential example of nondifferentiability. Indeed, gradient-based attacks are only effective when models are usefully differentiable. And yet, the paper states it verifies “the absence of gradient obfuscation”.
Figure 1: Buggy differentiable rounding operation.
And this is the reason why diff_round is not just a single function call to torch.round: the diff_round function is designed to implement a differentiable rounding operation. Unfortunately this function is not correct, and has a significant bug.
Figure 1 shows a plot of this diff_round function when varying the input from 0.0 to 0.3, and then a zoomed in view of the near-constant region. Observe that the gradient of the function is negative almost everywhere. And so, therefore, when we perform gradient descent it will actually move away from the adversarial examples!
And so while it is technically true that this function is “differentiable” in the sense that $\nabla f(x)$ is defined almost everywhere, Sabre contains the single most extreme form of gradient masking the author of this paper has ever seen. Not only does the gradient hide the true direction of the adversarial examples, it actually points directly away from them!
6.3 Our second break
Previously we broke Sabre by removing an unnecessary BPDA wrapper. We now insert a necessary one.
1diff --git a/core/defenses/sabre.py b/core/defenses/sabre.py
2index 319aebd..69fec01 100644
3--- a/core/defenses/sabre.py
4$+$ b/core/defenses/sabre.py
5@@ -61,7 +61,7 @@ class Sabre(nn.Module):
6 def precision_blend(self, x):
7 if self.eps > 0:
8 precision = max(min(-int(math.floor(math.log10(abs(1.25 * self.eps)))) - 1, 1), 0)
9- x = self.diff_round(x, decimals=precision)
10+ x = BPDAWrapper(lambda x: self.diff_round(x, decimals=precision))(x)
11 return x
On MNIST, where the defense binarizes inputs to either 0 or 1, we achieve an attack success rate of 21%. This is significantly higher than the 13% attack success rate of discretization alone, as reported by [8] which was first published in 2017. Thus Sabre with binarization is worse than binarization alone. In order to demonstrate conclusively that it is the binarization alone that has caused this robustness increase, if we adjust the number of decimals from 0 to 1 (and therefore round each input pixel to one of $\{0,0.1,0.2,\dots,0.9,1.0\}$), then we again see that our one-line diff again increases the attack success rate to 100%.
On CIFAR-10, where the defense sets the number of decimals to 1 and so again discretizes each input pixel to $\{0,0.1,0.2,\dots,0.9,1.0\}$, we achieve an attack success rate of 100%. Unfortunately, this CIFAR-10 result is for a model we have trained ourselves (but using the code published by the authors), and not for a model trained by the authors themselves. When we informed the authors of this second attack, they stated that this fix was incomplete for CIFAR-10, and that the defense must be adjusted differently for this model. However the authors have been unable to provide a description of the necessary changes, an implementation of the code for the fix, or a pointer to where the necessary adjustments are described in the original paper.
7 Author Response
We shared an advance copy of this paper describing the evaluation flaws from Sections 3-5 with the authors of Sabre. They provided the following response, which we quote verbatim:
{adjustwidth}
1.4cm1.8cm
7.1 Claims
  • Use of GPT-4: Generative models should not be used to judge research papers. Among the many known issues of such models are their tendency to produce factually incorrect information and having their responses biased by their inputs.
  • Attacking the model makes it more accurate: The proposed framework is primarily designed around the notion of noise reduction, such that the higher the amount of noise observed, the stronger the filtering. Therefore, the observation that attacking samples results in higher classification accuracies than those of benign samples is simply indicative of the presence of noise in the benign samples. Through the denoising of SABRE, noise is removed both from the attacked and benign samples, allowing classifiers to be less susceptible to noise, thereby also improving their generalization.
  • Nontrivial accuracy at $\varepsilon=0.5$: The evaluation results reported by the paper for an attack performed with $\varepsilon=0.5$ assume the attacks are performed on a model robustly trained using the SABRE defense. While the expectations set forth by Carlini et al. are theoretically sound, this does not guarantee that any attack method would find such perturbations.
7.2 Evaluation
The items below address the points made regarding the standard evaluation procedures highlighted in the critique:
  • (§5.2) “Verify iterative attacks perform better than single-step attacks” This would be the case if the iterative attacks only applied perturbations when the samples are correctly classified. This is however not the case with the baseline implementations typically used for benchmarking.
  • (§5.4) “Verify adaptive attacks perform better than any other” We considered “adaptive attacks” to be attacks that choose the most appropriate attack per sample given the defense, possibly out of multiple options. While the critique considers “adaptive attacks” to be attacks specifically designed against the defense, we believed designing such attacks to be outside the scope of the paper and left it for future work.
  • (§4.4) “Try at least one gradient-free attack and one hard-label attack” Gradient-free and hard-label attacks were not in the context of our threat model.
  • (§5.1) “Compare against prior work and explain important differences” The paper includes comparisons against several, widely used, prior work, with parameters set to allow fair comparison of different defenses within our evaluation settings. Unlike the reported robust accuracy of models that have been adversarially trained with a 40-step PGD attack, our evaluations aim to assess the robustness and generalization of models that have been adversarially trained with a 7-step PGD attack, which explains the drop in robustness accuracy.
  • (§4.11) “Attack with random noise of the correct norm” The threat model considered in the paper only considers gradient-based (whitebox) attacks.
  • (§4.13) “Perform ablation studies with combinations of defense components removed” Appendix A in the paper provides ablation studies of SABRE with and without non-core components of the proposed defense.
  • (§4.3) “Applying many nearly-identical attacks is not useful” The paper includes five distinct types of adversarial attacks widely used in the literature, namely: FGSM, PGD, EoT-PGD, CW, and AutoAttack.
7.3 SABRE Details
  • Evaluation against adaptive attacks. We considered “adaptive attacks” to be attacks that choose the most appropriate attack per sample given the defense, possibly out of multiple options. While the critique considers “adaptive attacks”. to be attacks specifically designed against the defense, we believed designing such attacks to be outside the scope of the paper and left it for future work.
  • Use of BPDA. BPDA was used for the simple purpose of ensuring differentiability regardless of the technical specificities of the implementation.
  • Comparison against a known-insecure defense The paper uses the ME-NET defense as a baseline because even though it has been broken, the approach is somewhat similar to SABRE and provides high levels of robustness when evaluated as proposed in the original paper. The paper references ME-Net for benchmarking and does not to suggest it as a secure standalone solution. The ME-Net defense was broken by the AutoAttack method, which has also been used to evaluate the robustness of SABRE.
  • Implementation of Adversarial Training. The paper uses a publicly available implementation of Adversarial Training, which is used in the literature for benchmarking. Unlike the reported robust accuracy of models that have been adversarially trained with a 40-step PGD attack, the SABRE paper assesses the robustness and generalization of models that have been adversarially trained with a 7-step PGD attack, explaining the drop in robustness accuracy.
7.4 Our response
We have several disagreements with the above response, but for brevity, limit our reply to two items:
  • On “adaptive attacks” and robustness to future attacks. As we write throughout this paper, the phrase “adaptive attack” has a standard definition in the adversarial machine learning literature: they are ones that are “adapted to the specific details of the defense and attempt to invalidate the robustness claims that are made” [3].
    Nowhere in the original paper do the authors state their (new) definition of adaptive attacks (running a set of fixed attacks and reporting the minimum among these), or state their proposed defense is only intended to be robust against the specific attacks considered in this paper. Moreover, it would be extremely abnormal for a security paper to ever make this claim because any defense, once published, will be subject to scrutiny. If robustness to fixed attacks is all that is desired, hundreds of papers already satisfy this goal.
  • On threat models. The authors claim gradient-free attacks (and attacks that just add noise of the correct norm) are not “in scope”. This is also incorrect: any attacker who can compute gradients can just throw away the gradient and only look at the loss. The threat model of white-box attacks encompasses the threat model of black-box attacks.
8 Conclusion
Sabre is the second defense to be accepted at IEEE S&P in as many years that had clear signs of a flawed evaluation which should have been easily identified by qualified reviewers. As we have shown, these evaluation flaws are caused by one (or two) bugs in the original evaluation that can be fixed by modifying one (or two) lines of code.
There is no doubt that defending against adversarial examples remains an interesting and important problem. And reviewers are attracted to papers making bold claims: it is much more tempting to accept a paper that claims near-perfect robust accuracy than a paper that claims a marginal improvement in accuracy. But it is of utmost importance that papers with errors such as this are identified during review, so as to not suggest that incorrectly evaluating defenses is tolerable.
Where do we go from here?
Adversarial machine learning is no longer a field that only studies toy problems to develop a better scientific understanding of the robustness of machine learning models. There are now production systems that depend on the robustness of the underlying machine learning models.
If, as a community, we can not tell the difference between a defense that claims 90%+ robustness and one that actually gives 0% robustness—on the simplest datasets and with glaring evaluation flaws—what hope do we have for evaluating the robustness of defenses designed to protect real, complex, production systems?

### v2

@@
 But this fix contains a second bug; modifying one more line of code reduces robust accuracy to below baseline levels.
+After we released the first version of our paper online, the authors introduced another change to the defense; by commenting out one line of code during attack we reduce the robust accuracy to 0% again.
 1 Introduction
@@
 Unfortunately, this modification contains a second bug, and by modifying a second line of code we again completely break the defense.
-The authors claim to have had a fix to this second attack for six weeks, but we have been unable to verify these claims because the authors have not published code for this fix, nor have they described the necessary changes.
-When the authors provide this information we will update our paper with an analysis of any future modifications or corrections to the currently published defense implementation.
+We then released the first version of our paper online.
+After this initial release of this paper, the authors updated the code again with another fix; but this contained a third bug, and by commenting out a single line code when running the attack (and putting it back at test time), we reduced defense accuracy to 0%.
 2 Background & Sabre
@@
 6 Our attack
-We break Sabre twice.
+We break Sabre three times.
 6.1 Our first break
@@
 Throughout the paper, Sabre claims to be “end-to-end differentiable to avoid gradient masking”, claims “innate differentiability of [their] defense mechanism”, and claims to “ensure end-to-end differentiability” [5].
-But with this new changed proposed by the authors, this is no longer the case: discretization is the quinticential example of nondifferentiability.
+But with this new changed proposed by the authors, this is no longer the case: discretization is the quintessential example of nondifferentiability.
 Indeed, gradient-based attacks are only effective when models are usefully differentiable.
@@
 When we informed the authors of this second attack, they stated that this fix was incomplete for CIFAR-10, and that the defense must be adjusted differently for this model.
-However the authors have been unable to provide a description of the necessary changes, an implementation of the code for the fix, or a pointer to where the necessary adjustments are described in the original paper.
+The following section will break this modified defense.
+6.4 Our third break
+The first version of this paper was released on May 6th 2024, and as described above, attacked a model we had trained ourselves because the authors were unable to provide a pretrained model.
+Our initial paper contained the disclaimer “When the authors provide this information we will update our paper with an analysis of any future modifications or corrections to the currently published defense implementation.” On May 24th and 27th, the authors made several corrections to the code, also released a pretrained CIFAR-10 model.
+On the same day (May 27th) we broke this third release of the code, and updated the arXiv version of this paper.
+The authors’ 300-LoC diff makes many changes to many pieces of the code, again introducing changes not described in the original paper.
+We omit a more complete discussion surrounding these changes from this paper, but only remark that this is the second time the defense has been dramatically altered since it was initially released by the authors.
+We firmly believe that code for research papers should accurately reflect the experiments actually performed when writing the paper, and should not be altered in major ways after the fact.
+But for the purpose of our paper here it is not necessary to address this topic, because only one of the changes matters: the denoising forward pass has been updated to the following method:
+1def forward(self, inputs, x0):
+2 b, _, m, n = inputs.shape
+3 c = self.in_channels
+4
+5 inputs = reshape_inputs(inputs)
+6 x0 = reshape_inputs(x0)
+7
+8 inputs = torch.cat((inputs, x0), dim=1)
+9
+10 inputs = ((inputs*(v := ((1<<3)-1))).round() - ((inputs<0) & (inputs%1 != 0)).float())/v
+11 outputs = self.layers(inputs)
+12
+13 if math.sqrt(m * n) > int(math.sqrt(m * n)):
+14 outputs = outputs.reshape(b, c, -1)[:, :, :m * n].reshape(b, c, m, n)
+15 return outputs
+Note that line 10 of this code (again) completely discretizes the inputs—this time without any attempted smoothing.
+This causes the gradient to be identically zero, and so PGD makes no progress during attack.
+As a result, the defense appears to obtain 90% robust accuracy on CIFAR-10.
+To break the defense, we simply comment out this line of code when running the attack, and then put the line back in when running the final evaluation.
+This is essentially equivalent to wrapping the code with a BPDA block, but even simpler: because the inputs are being discretized to one of eight different levels, it is sufficient to pretend (during the attack) that there is no discretization at all.
+Below is our complete break:
+1diff --git a/models/denoise.py b/models/denoise.py
+2index ba8154c..472fbec 100644
+3--- a/models/denoise.py
+4$+$ b/models/denoise.py
+5@@ -38,7 +38,7 @@ class DenoisingCNN(nn.Module):
+6 inputs = torch.cat((inputs, x0), dim=1)
+7
+8- inputs = ((inputs*(v := ((1<<3)-1))).round() - ((inputs<0) & (inputs%1 != 0)).float())/v
+9+ #inputs = ((inputs*(v := ((1<<3)-1))).round() - ((inputs<0) & (inputs%1 != 0)).float())/v
+10
+11 outputs = self.layers(inputs)
+With this single change, we reduce the defense accuracy on the paper’s official CIFAR-10 implementation with the provided CIFAR-10 model to $0\%$ robust accuracy.
+As can be seen from this attack, one must be exceptionally careful when introducing defense components that may cause gradient masking.
+Each of the three releases of this paper contained a different, but significant, gradient masking operation, despite the paper’s claims it “ensure[s] end-to-end differentiability”.
 7 Author Response
@@
 Sabre is the second defense to be accepted at IEEE S&P in as many years that had clear signs of a flawed evaluation which should have been easily identified by qualified reviewers.
-As we have shown, these evaluation flaws are caused by one (or two) bugs in the original evaluation that can be fixed by modifying one (or two) lines of code.
+As we have shown, these evaluation flaws are caused by between one and three bugs in the original evaluation that can be fixed by modifying one to three lines of code.
 There is no doubt that defending against adversarial examples remains an interesting and important problem.

### v3

@@
 After we released the first version of our paper online, the authors introduced another change to the defense; by commenting out one line of code during attack we reduce the robust accuracy to 0% again.
+Then, in response to this attack, the authors released another change to the code, reverting many changes and introducing new defense components—again; we break this version of the code by adding a single minus sign to one of the attack hyperparmaeters.
 1 Introduction
@@
 [2])
-In this short paper we show that by changing a single line of code to fix a bug in the published evaluation code, the attack success rate increases to 100%.
-The authors responded to our attack by modifying the published defense code to introduce a new defense component that was not present in the original paper, thereby changing the defense to a different defense than was initially proposed.
-Unfortunately, this modification contains a second bug, and by modifying a second line of code we again completely break the defense.
-We then released the first version of our paper online.
-After this initial release of this paper, the authors updated the code again with another fix; but this contained a third bug, and by commenting out a single line code when running the attack (and putting it back at test time), we reduced defense accuracy to 0%.
+In this short paper we show that SABRE’s impossible levels of accuracy can be explained by a bug in the original evaluation; fixing the bug by changing a single line of code increases the attack success rate increases to 100%.
+In response, in a process that will repeat three more times, the authors make a change to the code and introduce defense components never described in the original paper.
+We show that each of these three changes contains bugs that invalidate the results we break: the first by inserting a single function call; the second by commenting out a single line of code; and the third by negating one attack hyperparameter.
 2 Background & Sabre
@@
 Each of the three releases of this paper contained a different, but significant, gradient masking operation, despite the paper’s claims it “ensure[s] end-to-end differentiability”.
+6.5 Our fourth break
+On June 4th the authors released a fourth version of their code in response to our prior attack.
+The purpose of this change, taken from the git commit title, is “refactoring with no discretization”.
+That is, it it completely removes all discretization steps that were introduced in the second and third revisions of the code.
+[^1] It also makes many other changes, again without explanation for how these changes relate to content described in the paper.
+Because these other changes do not impact our results we omit any further discussion.
+When asked if this was intentional, the authors stated that it was; and they assert this latest version of the code accurately reflects the experiments performed in the original paper.
+At the same time, in prior communication the authors asserted on multiple occasions that the discretization was a necessary component of their defense that was unintentionally omitted from the first code release.
+Both of these can not be true simultaneously.
+The authors have now asserted that the code used in the experiments both did—and did not—include discretization.
+In this section we briefly study whether or not this version of the code could have been the code used for the original experiments in the paper.
+Then, we show that (regardless of whether or not this code was original) this implementation is also not robust to attack.
+Unreproducible.
+We find that this specific version of the code can not have been what was used to perform the original experiments: without any modifications, when we run the provided defense evaluation with the reference PGD attack on a CIFAR-10 model trained by the authors we find that it increases the defense accuracy to 100%.
+This is despite the original paper stating that running PGD on a CIFAR-10 trained model should yield 93% accuracy.
+(We are unable to run the MNIST model because the authors’ changes have since broken this model and its clean accuracy is now 10%.)
+Our attack.
+Nevertheless, we now break this version of the defense awaiting any future revisions.
+Our attack here is entirely trivial: whereas previously we have §6.1 removed a function call, §6.2 inserted a function call, and §6.3 commented out a line of code, in this section we now insert a single minus sign.
+Specifically, we make the following observation: if performing PGD stepping in the direction of the gradient increases accuracy when under attack—why shouldn’t we just step in the opposite direction of the gradient to decrease accuracy when under attack?
+To do this, all we need to do is negate the value of $\alpha$: the attack’s step size hyperparameter.
+1diff --git a/notebooks/cifar10.ipynb b/notebooks/cifar10.ipynb
+2index 543aa90..f1e5553 100644
+3--- a/notebooks/cifar10.ipynb
+4$+$ b/notebooks/cifar10.ipynb
+5@@ -389,7 +389,7 @@
+6 "metadata": {},
+7 "outputs": [],
+8 "source": [
+[…]
+[…]
+[…]
+[…]
+[…]
+We find this attack completely breaks the latest version of the code: negating the step direction reduces defense accuracy to 4%, well below the random guessing baseline.
+An explanation of this phenomenon is beyond the scope of this paper.
+It could be because the model was retrained (again) in a different way and this retraining caused gradients to point backwards; or it could be because the reconstruction model now behaves differently than it did previously.
+Given the simplicity of this attack, we aim to confirm it is effective by providing the authors with a list of adversarial examples we generated by negating the step size.
+Despite several attempts, the authors have been unwilling to verify the efficacy of these adversarial examples.
 7 Author Response